emilyiskewl:

What is the equation of a line passing through (-3, 7) and having a slope of -1/2 ?

Narad:

The equation of a line passing through the point (a,b) and having a slope of m is \[y-b=m(x-a)\]

emilyiskewl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Narad The equation of a line passing through the point (a,b) and having a slope of m is \[y-b=m(x-a)\] \(\color{#0cbb34}{\text{End of Quote}}\)